# Satisficing Regret Minimization in Bandits

### v2

@@ Appendix B Proof of Theorem 2
$\mathbb{E}\left[\sum_{s=1}^{t_{i}}\max\{S-r(X_{s}),0\}\right]\leq\mathbb{E}\left[t_{i}r(X^{*})-\sum_{s=1}^{t_{i}}r(X_{s})\right]\leq C_{1}t_{i}^{\alpha}\log(t_{i})^{\beta}\leq\dfrac{2C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}^{-\alpha/(1-\alpha)}\log(2/\gamma_{i})^{\beta}.$
@@
$\mathbb{E}[r(X^{*})-r(\hat{X}_{i})]\leq\dfrac{2C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}\log(2/\gamma_{i})^{\beta}.$
Therefore the satisficing regret incurred when pulling $\hat{X}_{i}$ is bounded by
@@
Since $r(\hat{X}_{i})\leq r(X^{*})<S$, for every $k$, round $i$ is not terminated with probability at most
@@
$\leq$ $\text{Pr}\left(\dfrac{\hat{r}^{\text{tot}}_{i}}{T_{i}+k}--\sqrt{\dfrac{4\log(T_{i}+k)}{T_{i}+k}}>r(\hat{X}_{i})\right)$
@@
Therefore expected number of arm pulls of $\hat{X}_{i}$ after pulling $T_{i}$ times is bounded by
@@
[…]
@@ C.1 Proof of Theorem 3
We denote $\texttt{Regret}_{S}^{(i)}$ as the satisficing regret under instance $i$. Then by the Bretagnolle-Huber inequality (see, e.g., Theorem 14.2 of Lattimore and Szepesvári (2020)), we have
$\texttt{Regret}_{S}^{(1)}+\texttt{Regret}_{S}^{(2)}\geq$ $\Delta\sum_{t=1}^{T_{0}}\text{Pr}_{t}^{(1)}(X_{t}=2)+\text{Pr}_{t}^{(2)}(X_{t}=1)$
@@ C.2 Proof of Theorem 4
Assume without loss of generality that $\Delta\leq 1$. Denote $T_{0}=\lceil 1/16\Delta^{2}\rceil$. We consider the following instances of two-armed bandit: $\mathcal{X}=[-3,3]$, and reward function is defined as follows:
@@
We denote $\texttt{Regret}_{S}^{(i)}$ as the satisficing regret under instance $i$, then by the Bretagnolle-Huber inequality (see, e.g., Theorem 14.2 of Lattimore and Szepesvári (2020)), we have
$\texttt{Regret}_{S}^{(1)}+\texttt{Regret}_{S}^{(2)}$ $\geq\Delta\sum_{t=1}^{T_{0}}\text{Pr}_{t}^{(1)}(X_{t}\geq 0)+\text{Pr}_{t}^{(2)}(X_{t}\leq 0)$

### v3

@@ Appendix B Proof of Theorem 2
$\mathbb{E}\left[t_{i}r(A^{*})-\sum_{s=1}^{t_{i}}r(A_{s})\right]\leq C_{1}t_{i}^{\alpha}\log(t_{i})^{\beta}\leq\dfrac{2C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}^{-\alpha/(1-\alpha)}\log(2/\gamma_{i})^{\beta}.$
@@
$\mathbb{E}[r(A^{*})-r(\hat{A}_{i})]\leq\dfrac{2C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}\log(2/\gamma_{i})^{\beta}.$
Therefore the satisficing regret incurred when pulling $\hat{A}_{i}$ is bounded by
@@
Since $r(\hat{A}_{i})\leq r(A^{*})<S$, for every $k$, round $i$ is not terminated with probability at most
@@
$\leq$ $\text{Pr}\left(\dfrac{\hat{r}^{\text{tot}}_{i}}{T_{i}+k}--\sqrt{\dfrac{4\log(T_{i}+k)}{T_{i}+k}}>r(\hat{A}_{i})\right)$
@@
Therefore expected number of arm pulls of $\hat{A}_{i}$ after pulling $T_{i}$ times is bounded by
@@
$\mathbb{E}[\texttt{Regret}]\leq$ $\sum_{i=1}^{i_{\max}}\dfrac{6C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}^{-\alpha/(1-\alpha)}\log(2T)^{\beta}$
@@ C.1 Proof of Theorem 3
We denote $\texttt{Regret}_{S}^{(i)}$ as the satisficing regret under instance $i$. Then by the Bretagnolle-Huber inequality (see, e.g., Theorem 14.2 of Lattimore and Szepesvári 2020), we have
$\texttt{Regret}_{S}^{(1)}+\texttt{Regret}_{S}^{(2)}\geq$ $\Delta\sum_{t=1}^{T_{0}}\text{Pr}_{t}^{(1)}(A_{t}=2)+\text{Pr}_{t}^{(2)}(A_{t}=1)$
@@ C.2 Proof of Theorem 4
Assume without loss of generality that $\Delta\leq 1$. Denote $T_{0}=\lceil 1/16\Delta^{2}\rceil$. We consider the following instances of two-armed bandit: $\mathcal{A}=[-3,3]$, and reward function is defined as follows:
@@
We denote $\texttt{Regret}_{S}^{(i)}$ as the satisficing regret under instance $i$, then by the Bretagnolle-Huber inequality (see, e.g., Theorem 14.2 of Lattimore and Szepesvári 2020), we have
$\texttt{Regret}_{S}^{(1)}+\texttt{Regret}_{S}^{(2)}$ $\geq\Delta\sum_{t=1}^{T_{0}}\text{Pr}_{t}^{(1)}(A_{t}\geq 0)+\text{Pr}_{t}^{(2)}(A_{t}\leq 0)$